# Preserving Task-Relevant Information Under Linear Concept Removal

### v1

@@ B.1 Comparing the projections for different levels of regularization
In this subsection, we investigate how the difference in performance between SPLICE, LEACE and SAL changes as a function of regularization. We use the Bias in Bios dataset and conduct the experiment as outlined in Section 4.1, but instead of selecting the $l_{2}$ regularization based on a validation set we fix the level or regularization. The results of this procedure are shown in Figure 5 and 6. As we lower the level of $l_{2}$ regularization (e.g. $\lambda=0.0001$) the difference between the methods becomes indistinghuishable from zero.
Figure 5: The difference between the SPLICE and LEACE projections on the Bias in Bios dataset for different levels of $l_{2}$ regularization. We show the difference (worst-group) accuracy of SPLICE minus the (worst-group) accuracy of LEACE. We re-train the last-layer after applying each projection. Points are based on the average over 3 seeds. The error bars reflect the 95% confidence interval.
Figure 6: The difference between the SPLICE and LEACE projections on the Bias in Bios dataset for different levels of $l_{2}$ regularization. We show the difference (worst-group) accuracy of SPLICE minus the (worst-group) accuracy of SAL. We re-train the last-layer after applying each projection. Points are based on the average over 3 seeds. The error bars reflect the 95% confidence interval.
@@ B.2 Removal of covariance for the Bias in Bios dataset for different projections
To quantify the extent to which $\bm{\Sigma}_{\bm{x},y_{\mathrm{prof}}}$ is preserved, we measure the ratio of the squared $l_{2}$ norm of the transformed covariance $\mathbf{P}\bm{\Sigma}_{\bm{x},y_{\mathrm{prof}}}$ after the projection $\mathbf{P}$ and original covariance $\bm{\Sigma}_{\bm{x},y_{\mathrm{prof}}}$. Figure 7 shows the effect of changing the conditional probability $p(y_{\mathrm{prof}}=a\mid z_{\mathrm{gender}}=a)$ on this ratio. For SPLICE, by design, $\bm{\Sigma}_{\bm{x},y_{\mathrm{prof}}}$ is preserved regardless of $p(y_{\mathrm{prof}}=a\mid z_{\mathrm{gender}}=a)$, and the ratio remains 1. As $p(y_{\mathrm{prof}}=a\mid z_{\mathrm{gender}}=a)$ increases, SAL and LEACE lead to a removal of $\bm{\Sigma}_{\bm{x},y_{\mathrm{prof}}}$. For instance, at $p(y_{\mathrm{prof}}=a\mid z_{\mathrm{gender}}=a)=0.9$, after LEACE and SAL the ratio between the transformed and original covariances become respectively 0.07 and 0.001.
Figure 7: The ratio $\frac{||\mathbf{P}\bm{\Sigma}_{\bm{x},y_{\mathrm{prof}}}||^{2}_{2}}{||\bm{\Sigma}_{\bm{x},y_{\mathrm{prof}}}||^{2}_{2}}$ as a function of the relationship between $y_{\mathrm{prof}},z_{\mathrm{gender}}$
@@ B.3 The mean difference between the original and transformed images for the CelebA dataset
To verify that the dynamics illustrated in Figure 4 hold across images, we measure the average difference between the original image before and after a projection. This is shown in Figure 8 for all combinations of $z_{\mathrm{smiling}}$ and $y_{\mathrm{glasses}}$. For individuals with glasses & not smiling, SPLICE heavily accentuates the glasses by (on average) making them darker. For individuals without glasses & smiling, SPLICE makes the area around the eyes lighter.
Figure 8: The mean difference between the original image and after an projection for every combination of $z_{\mathrm{smiling}},y_{\mathrm{glasses}}$.
@@ B.4 Applying the projection to multiple layers
Bias in Bios: we apply the projections (SAL, LEACE, SPLICE) to one of the 5 last layers of a BERT model. In this case, we do not re-train the subsequent layers. The accuracy after this procedure, per layer, is provided in Figure 9, and for worst-group accuracy in Figure 10. For later layers, similar to the results reported in Section 4.1, SPLICE outperforms the other projections when the conditional probability $p(y_{\mathrm{prof}}=a\mid z_{\mathrm{gender}}=a)$ increases.
Per layer, we report the $||\bm{\Sigma}_{\bm{x},z_{\mathrm{gender}}}||_{2}$ and $||\bm{\Sigma}_{\bm{x},y_{\mathrm{prof}}}||_{2}$ in respectively Table 3 and 3. In the earlier layers (7-10), both $||\bm{\Sigma}_{\bm{x},z_{\mathrm{gender}}}||_{2}$ and $||\bm{\Sigma}_{\bm{x},y_{\mathrm{prof}}}||_{2}$ are relatively low, indicating relatively little covariance between the embeddings and $z_{\mathrm{gender}},y_{\mathrm{prof}}$. As $||\bm{\Sigma}_{\bm{x},z_{\mathrm{gender}}}||_{2}$ and $||\bm{\Sigma}_{\bm{x},y_{\mathrm{prof}}}||_{2}$ increase in later layers (11-12), the difference between the projections also becomes clearer.
Table 2: The $||\bm{\Sigma}_{\bm{x},z_{\mathrm{gender}}}||_{2}$ for the biography dataset per layer
@@
Table 3: The $||\bm{\Sigma}_{\bm{x},y_{\mathrm{prof}}}||_{2}$ for the biography dataset per layer
@@
Profession dataset: When applying the projections, we start by the first layer in the sequence of layers where we alter the embeddings. After determining the projection for the embeddings at this layer, we subsequently determine it for the next, taking into account the projection of the previous layer. Table 4 shows the result of this procedure on Llama 2 7B. It remains the case that after SPLICE applying SPLICE to multiple layers, the model relies to a greater extent on factual information than when using SAL or LEACE.
@@
Winobias dataset: Similar to the Profession dataset, we apply the projections to embeddings of subsequent layers, taking into account the projection at the previous layer. As illustrated per Figure 11, we observe that the performance of SPLICE decreases as we apply it to more layers. Potentially, this is because of the (large) dimensionality of $\bm{\Sigma}_{\bm{x},y_{\mathrm{profession}}}\in\mathbb{R}^{d\times 40}$. This result highlights a potential limitation of our projection.
@@
B.5 Applying the projections to the full Bias in Bios dataset
@@
For completeness, we also show results for applying each projection to the complete Bias in Bios dataset, with all 28 professions, in Table 5. Here, we observe that when not re-training the last layer, both SPLICE and LEACE outperform SAL. This is presumably because both SPLICE and LEACE are better able to preserve the original embeddings compared to SAL. When re-training, all methods become statistically indistinghuishable in terms of performance, despite being trained with regularization (in contrast to the set-up discussed in Theorem 3.2). Potentially this is because the relationship between the 28 professions and gender is not as strong as in the setting considered in Section 4.1.
@@ C.1 Datasets
Profession dataset: consists of a total of 320 professions. We randomly select 219 professions for the training set, and use the remaining 101 for the test set. This is different from Limisiewicz et al. [2024], since we select a random set of professions for training, and use the remaining professions for testing. Each profession has an associated $z_{\mathrm{stereo}}$ and $y_{\mathrm{fact}}$ score, based on Bolukbasi et al. [2016]. Per profession, we create 11 prompts, based on the prompt templates in Table 7. This leads to a training set that consists of 2409 samples. For the test set, we subsample 950 prompts (from a total of $1$) such that the correlation between the $z_{\mathrm{stereo}}$ and $y_{\mathrm{fact}}$ score is 0. This is in order to disentangle the relationship between each score and our variable of interest $\mathrm{odds}_{he/she,i}$ as defined in equation 7.
@@ C.2 Details on models and training procedure
Llama models: we use the base model of Llama 2 7B, Llama 2 13B and Llama 3.1 8B. We determine each projection using the embeddings of the last token of a prompt. During test time, we apply the projection to each token, after the embeddings are normalized via the $\mathrm{RMSNorm}$ operation. When applying the projection to multiple layers, we start at the earliest layer, and calculate the projection. Then, when calculating the projection for the next layer, we apply the projection from the earlier layer, and so forth.

### v2

@@ B.1 Comparing the projections for different levels of regularization
In this subsection, we investigate how the difference in performance between SPLINCE, LEACE and SAL changes as a function of regularization. We use the Bias in Bios dataset and conduct the experiment as outlined in Section 4.1, but instead of selecting the $l_{2}$ regularization based on a validation set we fix the level or regularization. The results of this procedure are shown in Figure 5 and 6. As we lower the level of $l_{2}$ regularization (e.g. $\lambda=0.0001$) the difference between the methods becomes indistinghuishable from zero.
Figure 5: The difference between the SPLINCE and LEACE projections on the Bias in Bios dataset for different levels of $l_{2}$ regularization. We show the difference (worst-group) accuracy of SPLINCE minus the (worst-group) accuracy of LEACE. We re-train the last-layer after applying each projection. Points are based on the average over 3 seeds. The error bars reflect the 95% confidence interval.
Figure 6: The difference between the SPLINCE and SAL projections on the Bias in Bios dataset for different levels of $l_{2}$ regularization. We show the difference (worst-group) accuracy of SPLINCE minus the (worst-group) accuracy of SAL. We re-train the last-layer after applying each projection. Points are based on the average over 3 seeds. The error bars reflect the 95% confidence interval.
@@ B.2 Removal of covariance for the Bias in Bios dataset for different projections
To quantify the extent to which $\bm{\Sigma}_{\bm{x},y_{\mathrm{prof}}}$ is preserved, we measure the ratio of the squared $l_{2}$ norm of the transformed covariance $\mathbf{P}\bm{\Sigma}_{\bm{x},y_{\mathrm{prof}}}$ after the projection $\mathbf{P}$ and original covariance $\bm{\Sigma}_{\bm{x},y_{\mathrm{prof}}}$. Figure 7 shows the effect of changing the conditional probability $p(y_{\mathrm{prof}}=a\mid z_{\mathrm{gender}}=a)$ on this ratio. For SPLINCE, by design, $\bm{\Sigma}_{\bm{x},y_{\mathrm{prof}}}$ is preserved regardless of $p(y_{\mathrm{prof}}=a\mid z_{\mathrm{gender}}=a)$, and the ratio remains 1. As $p(y_{\mathrm{prof}}=a\mid z_{\mathrm{gender}}=a)$ increases, SAL and LEACE lead to a removal of $\bm{\Sigma}_{\bm{x},y_{\mathrm{prof}}}$. For instance, at $p(y_{\mathrm{prof}}=a\mid z_{\mathrm{gender}}=a)=0.9$, after LEACE and SAL the ratio between the transformed and original covariances become respectively 0.07 and 0.001.
[…]
@@ B.3 The mean difference between the original and transformed images for the CelebA dataset
To verify that the dynamics illustrated in Figure 4 hold across images, we measure the average difference between the original image before and after a projection. This is shown in Figure 8 for all combinations of $z_{\mathrm{smiling}}$ and $y_{\mathrm{glasses}}$. For individuals with glasses & not smiling, SPLINCE heavily accentuates the glasses by (on average) making them darker. For individuals without glasses & smiling, SPLINCE makes the area around the eyes lighter.
Figure 8: The mean difference between the original image and after a projection for every combination of $z_{\mathrm{smiling}},y_{\mathrm{glasses}}$.
@@ B.4 Applying the projection to multiple layers
Bias in Bios: we apply the projections (SAL, LEACE, SPLINCE) to one of the 5 last layers of a BERT model. In this case, we do not re-train the subsequent layers. The accuracy after this procedure, per layer, is provided in Figure 9, and for worst-group accuracy in Figure 10. For later layers, similar to the results reported in Section 4.1, SPLINCE outperforms the other projections when the conditional probability $p(y_{\mathrm{prof}}=a\mid z_{\mathrm{gender}}=a)$ increases.
Per layer, we report the $\bm{\Sigma}_{\bm{x},z_{\mathrm{gender}}}||_{2}$ and $\bm{\Sigma}_{\bm{x},y_{\mathrm{prof}}}||_{2}$ in respectively Table 3 and 3. In the earlier layers (7-10), both $\bm{\Sigma}_{\bm{x},z_{\mathrm{gender}}}||_{2}$ and $\bm{\Sigma}_{\bm{x},y_{\mathrm{prof}}}||_{2}$ are relatively low, indicating relatively little covariance between the embeddings and $z_{\mathrm{gender}},y_{\mathrm{prof}}$. As $\bm{\Sigma}_{\bm{x},z_{\mathrm{gender}}}||_{2}$ and $\bm{\Sigma}_{\bm{x},y_{\mathrm{prof}}}||_{2}$ increase in later layers (11-12), the difference between the projections also becomes clearer.
Table 2: The $\bm{\Sigma}_{\bm{x},z_{\mathrm{gender}}}||_{2}$ for the biography dataset per layer
@@
Table 3: The $\bm{\Sigma}_{\bm{x},y_{\mathrm{prof}}}||_{2}$ for the biography dataset per layer
@@
Profession dataset: When applying the projections, we start by the first layer in the sequence of layers where we alter the embeddings. After determining the projection for the embeddings at this layer, we subsequently determine it for the next, taking into account the projection of the previous layer. Table 4 shows the result of this procedure on Llama 2 7B. It remains the case that after SPLINCE applying SPLINCE to multiple layers, the model relies to a greater extent on factual information than when using SAL or LEACE.
@@
Winobias dataset: Similar to the Profession dataset, we apply the projections to embeddings of subsequent layers, taking into account the projection at the previous layer. As illustrated per Figure 11, we observe that the performance of SPLINCE decreases as we apply it to more layers. Potentially, this is because of the (large) dimensionality of $\bm{\Sigma}_{\bm{x},y_{\mathrm{profession}}}\in\mathbb{R}^{d\times 40}$. This result highlights a potential limitation of our projection.
@@
B.5 Additional results for vision datasets
In this section, we briefly investigate the performance of SPLINCE and other projections on two vision datasets: Waterbirds Sagawa et al. [2020] and the CelebA dataset described in Section 4.3. Similar to the experiments in Section 4.1, we alter the extent to which the main-task co-occurs with the concept. For the Waterbirds, we seek to predict whether or not a land or waterbird is present ($y_{\mathrm{bird}}\in\{0,1\}$), while removing the concept of the land or water background, denoted $z_{\mathrm{back}}\in\{0,1\}$. For the CelebA dataset, we alter the extent to which an image with glasses $y_{\mathrm{glasses}}$ co-occurs with $z_{\mathrm{smiling}}$. Details on the datasets and training procedure can be found in Appendix C. We also compare SPLINCE to two benchmark out-of-distribution (OOD) generalization methods. First, deep feature reweighting (DFR, Kirichenko et al. [2022]), where a model is trained on a subsampled dataset, where each combination of the main-task and concept occurs with equal probability. We implement the version of DFR where the subsampled dataset comes from the training data, referred to as $\mathrm{DFR}_{\mathrm{TR}}$ in Kirichenko et al. [2022]. Second, we apply group distributional robust optimization (GDRO, Sagawa et al. [2020]) to the last layer. Details on the implementation of both methods can be found in Appendix C.2. We compare SPLINCE to these two methods for the vision datasets, as well as the NLP classification tasks outlined in 4.1.
Figure 12 compares each projection for the Waterbirds and CelebA dataset. For the Waterbirds dataset, the performance of each projection ( SAL, LEACE, SPLINCE) strongly deteriotates as the correlation between the main-task and the concept increases. For the CelebA dataset, ERM gives a superior performance compared to the projections. These results indicate that concept-removal methods such as SPLINCE, as well as SAL and LEACE, perform relatively worse on the last-layer embeddings of vision datasets rather than those generated for NLP tasks. This is in line with previous work [Holstege et al., 2024] and an interesting direction for future research.
Figure 12: Performance of different projections on the Waterbirds and CelebA dataset. We re-train the last-layer after applying each projection. Points are based on the average over 5 seeds for each of the two datasets. The error bards reflect the 95% confidence interval.
In Figure 13 we present the comparison of SPLINCE to $\mathrm{DFR}_{\mathrm{TR}}$ and GDRO for the Waterbirds and CelebA dataset. Both methods strongly outperform SPLINCE. It is worth emphasizing that SPLINCE is not explicitly designed to achieve strong out-of-distribution (OOD) generalization - rather to achieve certain fairness guarantees (e.g. linear guardedness) while maintaining main-task performance. In Figure 13 we present the comparison of SPLINCE to $\mathrm{DFR}_{\mathrm{TR}}$ and GDRO for the Bias in Bios and Multilingual Text Detoxification dataset. SPLINCE performs similar to both methods for the Bias in Bios dataset, and outperforms both (at a high correlation) for the Multilingual Text Detoxification dataset. This result is further empirical evidence that SPLINCE performs better for NLP tasks than vision datasets - as it is able to perform similar or better than methods designed for OOD generalization ( $\mathrm{DFR}_{\mathrm{TR}}$ and GDRO) for these two datasets. One potential reason that SPLINCE strongly outperforms $\mathrm{DFR}_{\mathrm{TR}}$ and GDRO for the Multilingual Text Detoxification dataset is that there is a greater number of possible combinations of the main-task and concept, as well as a smaller sample size, causing $\mathrm{DFR}_{\mathrm{TR}}$ and GDRO to overfit on the training data.
Figure 13: Performance of SPLINCE compared to deep feature reweighting ($\mathrm{DFR}_{\mathrm{TR}}$) and Group Distributional Robust Optimization (GDRO) on the Waterbirds and CelebA dataset. Each method is applied to the last layer embeddings. Points are based on the average over 5 seeds for each of the two datasets. The error bards reflect the 95% confidence interval.
Figure 14: Performance of SPLINCE compared to deep feature reweighting ($\mathrm{DFR}_{\mathrm{TR}}$) and Group Distributional Robust Optimization (GDRO) on the Bias in Bios and Multilingual Text Detoxification dataset. Each method is applied to the last layer embeddings. Points are based on the average over 5 seeds for each of the two datasets. The error bards reflect the 95% confidence interval.
B.6 Additional results for Large Language Models
In this section, we investigate SPLINCE and other projections on the same language tasks as outlined in 4.2, but for two additional LLMs: the Mistral v0.3 7B model [Jiang et al., 2023], and the Phi-2 model [Javaheripi et al., 2023] from Microsoft. In both cases, we use the base models as available on Huggingface, and we apply each projection to the last layer embeddings.
The results for the profession dataset are presented in Table 5, and for the Winobias dataset in Figure 15. For the profession dataset, the results are in line with the results for the Llama models as presented in Table 1. After applying each projection, the extent to which the models rely on factual information is greatly reduced, but this reduction is smallest for SPLINCE. For the Winobias dataset, we observe little to no change for the Phi-2 model after any of the projections. Most likely this is related to the overall poor performance of the Phi-2 model on this task, potentially due to its relatively smaller size compared to the other models (2.7B parameters). For the Mistral 7B v0.3 model, we observe an increase in the accuracy on anti-stereotypical prompts after applying SPLINCE.
Table 5: Results of applying different projections to the last layer of the Mistral 7B v0.3 and Phi-2 models for the profession dataset.
Mistral 7B v.03 Original 3,70 24,62
+SAL 0,95 4,86
+LEACE 1,34 10,0
+SPLINCE 1,37 14,76
Phi-2 (2.7B) Original 1,77 15,72
+SAL 0,77 5,44
+LEACE 0,77 7,58
+SPLINCE 0,74 11,23
Figure 15: Results of applying different projections to the last layer of the Mistral 7B v0.3 and Phi-2 models for the Winobias dataset. The left plot shows The left plot shows the accuracy on a test set consisting of half pro-stereotypical and half anti-stereotypical prompts. The right plot shows the accuracy on the anti-stereotypical prompts in this test set.
B.7 Applying the projections to the full Bias in Bios dataset
@@
For completeness, we also show results for applying each projection to the complete Bias in Bios dataset, with all 28 professions, in Table 6. Here, we observe that when not re-training the last layer, both SPLINCE and LEACE outperform SAL. This is presumably because both SPLINCE and LEACE are better able to preserve the original embeddings compared to SAL. When re-training, all methods become statistically indistinghuishable in terms of performance, despite being trained with regularization (in contrast to the set-up discussed in Theorem 3.2). Potentially this is because the relationship between the 28 professions and gender is not as strong as in the setting considered in Section 4.1.
@@ C.1 Datasets
Profession dataset: consists of a total of 320 professions. We randomly select 219 professions for the training set, and use the remaining 101 for the test set. This is different from Limisiewicz et al. [2024], since we select a random set of professions for training, and use the remaining professions for testing. Each profession has an associated $z_{\mathrm{stereo}}$ and $y_{\mathrm{fact}}$ score, based on Bolukbasi et al. [2016]. Per profession, we create 11 prompts, based on the prompt templates in Table 8. This leads to a training set that consists of 2409 samples. For the test set, we subsample 950 prompts (from a total of 1111) such that the correlation between the $z_{\mathrm{stereo}}$ and $y_{\mathrm{fact}}$ score is 0. This is in order to disentangle the relationship between each score and our variable of interest $\mathrm{odds}_{he/she,i}$ as defined in equation 7.
@@
Waterbirds dataset: introduced by Sagawa et al. [2020], it is a combination of the Places dataset [Zhou et al., 2016] and the CUB dataset [Welinder et al., 2010]. A ‘water background’ is set by selecting an image from the lake and ocean categories in the places dataset, and the ‘land background’ is set based on the broadleaf and bamboo forest categories. A waterbird/land is then pasted in front of the background. When creating new versions of the dataset, we change the $p(y_{\mathrm{bird}}=a\mid z_{\mathrm{back}}=a)$ for $a\in{0,1}$ and keep the size of the training set at 4,775/1,199 for the training and validation set respectively. For the test set, we select 5,796 samples where $p(y_{\mathrm{bird}}=a\mid z_{\mathrm{back}}=a)=0.5$.
@@ C.2 Details on models and training procedure
Llama models: we use the base model of Llama 2 7B, Llama 2 13B and Llama 3.1 8B as available on Huggingface. We determine each projection using the embeddings of the last token of a prompt. During test time, we apply the projection to each token, after the embeddings are normalized via the $\mathrm{RMSNorm}$ operation. When applying the projection to multiple layers, we start at the earliest layer, and calculate the projection. Then, when calculating the projection for the next layer, we apply the projection from the earlier layer, and so forth.
@@
When implementing $\mathrm{DFR}_{\mathrm{TR}}$, we use a subsampled set from the training dataset where each group has an equal size. Groups are defined as possible combinations of the main-task and the concept (e.g. in the Waterbirds dataset, there are four groups, as $y_{\mathrm{bird}}\in\{0,1\}$ and $z_{\mathrm{back}}\in\{0,1\}$). For GDRO, we use a learning rate $\eta=0.1$ to update the weights per group after each gradient descent step, similar to Sagawa et al. [2020].
Vision datasets: For the Waterbirds dataset, we use the ResNet50 architecture implemented in the torchvision package: torchvision.models.ResNet50(pretrained=True). We finetune the model using the parameters of Kirichenko et al. [2022]: a learning rate of $10^{-3}$, a weight decay of $10^{-3}$, a batch size of 32, and for 100 epochs without early stopping. We use stochastic gradient descent (SGD) with a momentum parameter of $0.9$. For CelebA, we simply run a logistic regression, akin to last-layer retraining, on the downscaled grey-scale images.
Appendix D Ethical considerations
As with any technique that aims to ensure that the predictions of a machine learning (ML) model are fair, practitioners should exercise caution when deploying SPLINCE in real-world settings where decisions can affect people’s lives. Naturally, our work considers specific technical notions of fairness, and is evaluated on a limited number of datasets, that do not reflect all the considerations one should take into account in deployment. “Fairness" is a multifaceted construct, and our approach addresses only certain dimensions. Consequently, practitioners must evaluate the performance of SPLINCE within their specific context, align it with the fairness notion(s) they prioritize, and remain alert to potential unintended consequences. Importantly, SPLINCE targets a very specific definition of bias, quantified by the ability of a linear model to predict a protected attribute. The method is not necessarily expected to work for non-linear models, or for other definitions of fairness.